OpenStudy (anonymous):

simplify the 4th root of 80

OpenStudy (anonymous):

\[\sqrt[4]{80}\] \[\sqrt[4]{16 \times 5}\] Think you can continue from here?

OpenStudy (anonymous):

\[\sqrt{a \times b} = \sqrt{a} \times \sqrt{b}\] This rule might help.